Demand and supply equations for housing market per month are given below.

β€’ Demand for housing: 𝑄𝐷 = 2500 βˆ’ 0.5 𝑃

β€’ Supply of housing: 𝑄𝑆 = βˆ’500 + 𝑃

a) Draw the demand and supply curves for housing market in one graph.

b) Find the equilibrium quantity and price for housing.

c) Compute the consumer and producer surplus in equilibrium.Β 

d) Suppose that the government set a rent ceiling at $1800. What are the quantities ofΒ 

housing supplied and demanded at this price? In this case, is there a shortage orΒ 

surplus of houses?

e) How does the price ceiling affect the efficiency in the housing market?

f) Calculate the deadweight loss in the housing market after the price ceiling isΒ 

imposed by the government.

g) Calculate the potential spending for housing search activities


1
Expert's answer
2021-12-13T16:53:47-0500

a)



b) 𝑄𝐷 =𝑄𝑆

𝑄𝐷 = 2500 βˆ’ 0.5 𝑃

𝑄𝑆 = βˆ’500 + 𝑃

2500 βˆ’ 0.5 𝑃=βˆ’500 + 𝑃

3000=1.5P

P=2000

Q=1500

c)

"consumer surplus =\frac{(5000-2000)1500}{2}=1875000"


"producer surplus =\frac{(2000-500)1500}{2}=1125000"


d) P=1800

Q=1600

e)depending on the prices, there is a revival or a fading of the market

f)

"the deadweight loss=0.5(1800-2000)(1500-1600)=10000"

g) "(4000-1000)500=1500000"